Diabetes Screening

  1. Shown below are the real-life diagnostic characteristics of a particular diabetes-screening test – a fasting venous glucose measurement – from a hypothetical cross-sectional study population. Calculate the sensitivity, specificity and positive predictive value (PPV) of this screening test and estimate the prevalence of diabetes in the study population.

Table: Results of fasting venous glucose screening for diabetes.

Diabetes Screening Test

Actual diabetes status

Total

Present

Absent

Positive (+)

True +

67

False +

38

Total +

105

Negative (-)

False –

53

True –

1842

Total –

1895

Total diabetics

120

Total non-diabetics

1880

Total screened population

2000

Sensitivity:

Specificity:

PPV:

Prevalence of diabetes:

Selected Formulas:

Age-specific death rate = # of deaths per age stratum

Population per age stratum

Accident-specific death rate = # of deaths from accidents per age stratum

                                         Population per age stratum

Accident case-fatality rate = # of deaths from accidents per age stratum

Total # of accidents per age stratum

Mortality Rate = __# of deaths___

Population at Risk

Proportionate Mortality Ratio = # of deaths from accidents per age stratum

Total # of deaths per age stratum
Answer #1

Diabetes Screening Test

Actual diabetes status

Total

Present

Absent

Positive (+)

True +

67 (A)

False +

38 (B)

Total +

105

Negative (-)

False –

53 (C)

True –

1842 (D)

Total –

1895

Total diabetics

120

Total non-diabetics

1880

Total screened population

2000

Sensitivity = A/(A+C) × 100 = (67 / 120) x 100 = 55.83%

Specificity = D/(D+B) × 100 = (1842 / 1880) x 100 = 97.98%

PPV = A/(A+B) × 100 = (67 / 105) x 100 = 63.81%

Prevalence of diabetes = (Total diabetics / Total screened population) × 100 = (120 / 2000) x 100 = 6%